OpenStudy (beccab003):

Help with quick question. Medal to best answer. Thanks! f(x)=-2x-1 for all real numbers. What is f^-1(x), the inverse of f(x)?

OpenStudy (deadpoolroxz):

this link will give u the BEST ansewr http://www.webmath.com/nsolve.html

OpenStudy (beccab003):

I did it and it wouldn't go through. it said something about an incorrect variable. What variable do I put in?

OpenStudy (beccab003):

@campbell_st

OpenStudy (deadpoolroxz):

both

OpenStudy (campbell_st):

ok... to find the inverse swap x and y so if y = -2x - 1 make the swap x = -2y - 1 now to find the inverse, make y the subject of the equation.

OpenStudy (beccab003):

subject meaning find y?

OpenStudy (campbell_st):

yes, get a single y equal to something...

OpenStudy (beccab003):

so i plug in a number for y

OpenStudy (beccab003):

or x

OpenStudy (campbell_st):

no.... making y the subject add 1 to both sides of the equation. then divide both sides of the equation by -2

OpenStudy (beccab003):

so...adding 1 to both sides: 1x=-2y+0

OpenStudy (beccab003):

then dividing by -2= ???

OpenStudy (phi):

x = -2y - 1 ***so...adding 1 to both sides: 1x=-2y+0**** to add, write +1 on both sides: x+1 = -2y -1 + 1 you can simplify the right side (-1+1 is 0 and -2y + 0 is -2y) you get x+1 = -2y now divide both sides by -2 (that means divide *everything* on the left side by -2

OpenStudy (beccab003):

so 1/-2 = 0.5 so -2x+0.5 = -2y

OpenStudy (beccab003):

@phi

OpenStudy (phi):

the idea is you have an equation (two sides are equal) you divide each side "in half" and they are still equal to show "divide in half" we write /(-2) under the entire side. we do that to both sides can you do that for x+1 = -2y

OpenStudy (phi):

in other words draw a line under x+1, and then write -2 under the line on the right side, draw a line under -2y, then write -2 under that line can you do that ?

OpenStudy (beccab003):

yes I did and i figured out the answer. THanks so much